# Improving Offline-to-Online Reinforcement Learning with Q-Ensembles

### v3

@@ Appendix B Environment Settings
We investigate three MuJoCo locomotion tasks, namely HalfCheetah, Walker2d, and Hopper (Todorov et al., 2012). The goal of each task is to move forward as fast as possible, while keeping the control cost minimal. For each task, we consider four types of datasets. The random datasets consist of policy rollouts generated by random policies. The medium datasets contain rollouts from medium-level policies. The medium-replay datasets encompass all samples collected during the training of a medium-level agent from scratch. In the case of the medium-expert datasets, half of the data comprises rollouts from medium-level policies, while the other half consists of rollouts from expert-level policies. In this study, we exclude the random and the expert datasets, as in typical real-world scenarios, we rarely use a random policy or have an expert policy for system control. We utilize the v2 version of each dataset.
@@
We investigate the Antmaze navigation tasks that involve controlling an 8-DoF ant quadruped robot to navigate through mazes and reach a desired goal. The agent receives sparse rewards of +1/0 based on whether it successfully reaches the goal or not. We study each method using the following datasets from D4RL (Fu et al., 2020): large-diverse, large-play, medium-diverse, medium-play, umaze-diverse, and umaze. The difference between diverse and play datasets is the optimality of the trajectories they contain. The diverse datasets consist of trajectories directed towards random goals from random starting points, whereas the play datasets comprise trajectories directed towards specific locations that may not necessarily correspond to the goal. We use the v1 version of each dataset.
@@ Appendix C Experiment Details
For CQL, SAC and AWAC, we use the implementation provided by (Seno & Imai, 2022): https://github.com/takuseno/d3rlpy with default hyperparameters. For CQL-N and SAC-N, we keep the default setting from the CQL and SAC experiments other than the ensemble size N. For Balanced Replay (BR), as the official implementation provided by the author of (Lee et al., 2022) does not contain the offline pre-training part, we implement Balanced Replay based on d3rlpy. For LAPO-N, we extend the official implementation provided by the author of (Chen et al., 2022): https://github.com/pcchenxi/LAPO-offlineRL to easily adjust the size of ensemble. For PEX and IQL, we use the original implementation provided by the author of (Zhang et al., 2023): https://github.com/Haichao-Zhang/PEX. While we do not utilize the code provided by the original paper for certain methods, some of the our results obtained using the employed code demonstrate superior performance compared to those reported in the original paper. Moreover, when compared to the results provided in the original paper, our proposed E2O framework consistently outperforms them. We list the hyperparameters for these methods in Table 1.
@@
Hyperparameters CQL-N SAC-N LAPO-N AWAC BR PEX IQL
policy learning rate 3e-5 3e-5 2e-4 3e-4 3e-5 3e-4 3e-4
critic learning rate 3e-4 3e-4 2e-4 3e-4 3e-4 3e-4 3e-4
alpha learning rate 1e-4 1e-4 - - 3e-4 - -
VAE learning rate - - 2e-4 - - - -
value learning rate - - - - - 3e-4 3e-4
ensemble size 10 10 10 2 5 2 2
batch size 256 256 512 1024 256 256 256
@@
For all experiments, we report the online fine-tuning performance over 250K timesteps with 5 seeds. Specifically, our framework loads all pre-trained networks, including the policy network, ensemble Q network and ensemble target Q network, while appending the necessary temperature hyperparameter for SAC to facilitate further fine-tuning. In the Antmaze environment, we maintain the same reward modification as the offline phase during training but keep the rewards unchanged during evaluation. To ensure a fair comparison with IQL and PEX, which utilize offline data, we also load LAPO and E2O-LAPO with offline data for online fine-tuning.
@@ Appendix D Additional Results
In this section, we provide more experiments and detailed results to help understand our proposed E2O framework more comprehensively.
@@ D.1 Ablation on Offline Data
We conduct an ablation study to investigate the impact of using offline data during the online fine-tuning phase for all MuJoCo locomotion tasks, as shown in Figure 7. Our results show that E2O-CQL_buffer, which initializes the online buffer with offline data, exhibits slow performance improvement, while discarding the offline data allows it to achieve higher sample efficiency. This suggests that although many offline-to-online RL methods utilize offline data to alleviate performance degradation, it can adversely affect their online sample efficiency. In contrast, E2O-CQL successfully avoids significant performance drop even without using offline data, thereby enhancing learning efficiency during the online stage.
D.2 Comparison of LAPO and E2O-LAPO
Figure 8: Online learning curves of LAPO and E2O-LAPO across five seeds on Antmaze tasks. The solid lines and shaded regions represent mean and standard deviation, respectively.
On the Antmaze tasks, we have conducted a comparative analysis between our E2O-LAPO method and several offline-to-online RL methods. To further validate the effectiveness of our E2O framework, we directly utilize LAPO with two Q networks for offline-to-online training and compare it with our E2O-LAPO method. The results are shown in Figure 8. As original LAPO can achieve near-optimal performance in simple environments such as umaze and medium mazes during the offline stage, the online fine-tuning performance of both LAPO and E2O-LAPO is comparable. However, in the more challenging large maze environment, directly using the offline pre-trained LAPO agent for online fine-tuning leads to slow performance improvement. By employing our proposed E2O framework, we demonstrate that E2O-LAPO can not only enhance the offline performance of LAPO, but also facilitate more rapid performance improvement while maintaining the offline performance without degradation. This approach enables the offline agent to quickly adapt to the real-world environment, providing efficient and effective online fine-tuning.
@@ D.3 Visualization and Analysis
Figure 9: Visualization of the distribution of states generated by E2O-CQL, CQL in online phase and states in offline dataset.
To better understand the training efficiency of E2O in comparison to traditional pessimistic offline RL algorithms, we compare the distribution of states generated by CQL, E2O-CQL in the online phase, and the distribution of states from the offline dataset. To visualize the results clearly, we plot the distribution with t-Distributed Stochastic Neighbor Embedding (t-SNE) (Hinton & Roweis, 2002). The results are shown in Figure 9, it can be found that both the distribution of E2O-CQL states and CQL states bear some similarity to the distribution of offline states. However, the online states accessed by CQL are located on the edge of the offline area, but most still overlap with the offline states. On the other hand, the online states accessed by E2O-CQL deviate further from the offline states. With our ensemble-based design for optimistic exploration, E2O empowers the offline agent to explore more states beyond those contained in the offline dataset. This capability allows for swift adaptation to online environments and facilitates rapid performance improvement.
D.4 Detailed Results of E2O Components
In this section, we provide all learning curves of E2O components that are restricted by the length of the text in the main paper.
@@ D.4 Detailed Results of E2O Components
Figure 10 illustrates the performance of various offline-to-online RL approaches on MuJoCo locomotion tasks. It is evident that the OfflineRL normal-$\to$ OnlineRL method exhibits the best performance in the HalfCheetah environment. However, it demonstrates unstable learning in the more complex environments of Walker2d and Hopper. On the other hand, the OfflineRL normal-$\to$ OfflineRL approach remains stable but shows slower asymptotic performance. In contrast, the OfflineRL-N normal-$\to$ OnlineRL-N method no longer experiences performance collapse after transitioning to online fine-tuning, and its training process is relatively stable across all tasks. Additionally, OfflineRL-N normal-$\to$ OnlineRL-N achieves superior fine-tuned performance compared to OfflineRL normal-$\to$ OfflineRL. It is worth noting that the Hopper-medium-expert-v2 dataset represents a special case where all considered offline-to-online methods exhibit varying degrees of performance drop, as depicted in this figure and subsequent figures. Nevertheless, our E2O framework consistently achieves state-of-the-art performance in comparison to all baseline methods across most tasks.
@@
Figure 11 displays the performance of OnlineRL-N utilizing different Q-target computation methods on MuJoCo locomotion tasks. It is evident that MinQ exhibits remarkable stability across all tasks, albeit with slower performance improvement in the HalfCheetah and Walker2d environments. MeanQ and REM demonstrate excellent performance in the HalfCheetah environment, but struggle to improve in the more challenging environments of Walker2d and Hopper, and their learning process is characterized by instability. In contrast, RandomMinPair and WeightedMinPair showcase superior performance across most tasks, with the exception of the Hopper-medium-replay-v2 dataset where they exhibit slight instability in learning. Among these two methods, WeightedMinPair demonstrates slightly better stability and performance, thus we select it as the component of our final E2O framework and present the experiments related to RandomMinPair in the appendix.
@@
Figure 12 and Figure 13 present the performance of OnlineRL-N + WeightedMinPair and OnlineRL-N + RandomMinPair using different exploration methods on MuJoCo locomotion tasks, respectively. These two figures exhibit similar observations. In the Hopper environment, OAC achieves the best performance, but its performance improvement in HalfCheetah and Walker2d is relatively slow. The use of Bootstrapped DQN leads to minimal improvement in performance, while SUNRISE enhances the learning efficiency of OnlineRL-N + WeightedMinPair across most tasks, with the exception of the Hopper-medium-replay-v2 dataset where they exhibit slight instability in learning.
@@ D.5 Offline Performance
Figure 14 shows the offline performance of different methods on MuJoCo locomotion tasks. We observe that on certain datasets such as HalfCheetah-medium-v2 and HalfCheetah-medium-replay-v2, E2O-CQL exhibits a slight performance improvement compared to CQL. This indicates that the introduction of Q-ensembles indeed has some benefits for the performance in the offline stage. However, it is important to note that the use of Q-ensembles can impact the convergence speed. For instance, on the Hopper-medium-expert-v2 dataset, E2O-CQL demonstrates a noticeably slower convergence speed compared to CQL. Nevertheless, if both algorithms are allowed to continue training, for example, for 3M training steps, we believe that the performance of E2O-CQL can still surpass that of CQL.
Although different offline RL algorithms may have varying final performance in the offline stage, with some methods potentially performing worse than E2O-CQL, our comparisons in the online stage remain fair. On one hand, the introduction of Q-ensembles can enhance the performance of existing algorithms in the offline stage, which is a inherent advantage of Q-ensembles. On the other hand, higher performance in the offline stage can actually lead to performance drop in the online stage, while lower performance in the offline stage is less prone to such drop. Our proposed E2O framework ensures that the offline policy maintains high performance from the offline stage, and achieves rapid performance improvement in the online stage without encountering performance drop.
@@
Figure 15 displays the offline performance of various methods on Antmaze navigation tasks. Firstly, we observe that LAPO outperforms IQL in terms of offline performance, providing a higher starting point for the online phase. This is particularly evident in the umaze and medium maze environments, where LAPO nearly reaches the performance ceiling. Regarding LAPO and E2O-LAPO, since LAPO achieves near-optimal performance in simple environments such as umaze and medium mazes, their offline performance is comparable. However, in the more challenging large maze environment, the inclusion of Q-ensembles enables E2O-LAPO to surpass LAPO in terms of performance.
@@
ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v4

@@ Appendix B Environment Settings
We investigate three MuJoCo locomotion tasks, namely HalfCheetah, Walker2d, and Hopper Todorov et al. [2012]. The goal of each task is to move forward as fast as possible, while keeping the control cost minimal. For each task, we consider four types of datasets. The random datasets consist of policy rollouts generated by random policies. The medium datasets contain rollouts from medium-level policies. The medium-replay datasets encompass all samples collected during the training of a medium-level agent from scratch. In the case of the medium-expert datasets, half of the data comprises rollouts from medium-level policies, while the other half consists of rollouts from expert-level policies. In this study, we exclude the random and the expert datasets, as in typical real-world scenarios, we rarely use a random policy or have an expert policy for system control. We utilize the v2 version of each dataset.
@@
We investigate the Antmaze navigation tasks that involve controlling an 8-DoF ant quadruped robot to navigate through mazes and reach a desired goal. The agent receives sparse rewards of +1/0 based on whether it successfully reaches the goal or not. We study each method using the following datasets from D4RL Fu et al. [2020]: large-diverse, large-play, medium-diverse, medium-play, umaze-diverse, and umaze. The difference between diverse and play datasets is the optimality of the trajectories they contain. The diverse datasets consist of trajectories directed towards random goals from random starting points, whereas the play datasets comprise trajectories directed towards specific locations that may not necessarily correspond to the goal. We use the v1 version of each dataset.
@@ Appendix C Experiment Details
For CQL, SAC and AWAC, we use the implementation provided by Seno and Imai [2022]: https://github.com/takuseno/d3rlpy with default hyperparameters. For CQL-N and SAC-N, we keep the default setting from the CQL and SAC experiments other than the ensemble size N. For Balanced Replay (BR), as the official implementation provided by the author of Lee et al. [2022] does not contain the offline pre-training part, we implement BR based on d3rlpy. For LAPO-N, we extend the official implementation provided by the author of Chen et al. [2022]: https://github.com/pcchenxi/LAPO-offlineRL to easily adjust the size of ensemble. For PEX and IQL, we use the original implementation provided by the author of Zhang et al. [2023]: https://github.com/Haichao-Zhang/PEX. For Cal-QL, we use the implementation provided by Tarasov et al. [2022]: https://github.com/tinkoff-ai/CORL with default hyperparameters. While we do not utilize the code provided by the original paper for certain methods, some of the our results obtained using the employed code demonstrate superior performance compared to those reported in the original paper. Moreover, when compared to the results provided in the original paper, our proposed ENOTO framework consistently outperforms them. We list the hyperparameters for these methods in Table LABEL:Table:hyper.
@@
Hyperparameters CQL-N SAC-N LAPO-N AWAC BR Cal-QL PEX IQL
policy learning rate 3e-5 3e-5 2e-4 3e-4 3e-5 1e-4 3e-4 3e-4
critic learning rate 3e-4 3e-4 2e-4 3e-4 3e-4 3e-4 3e-4 3e-4
alpha learning rate 1e-4 1e-4 - - 3e-4 5e-3 - -
VAE learning rate - - 2e-4 - - - - -
value learning rate - - - - - - 3e-4 3e-4
ensemble size 10 10 10 2 5 2 2 2
batch size 256 256 512 1024 256 256 256 256
@@
For all experiments, we report the online fine-tuning performance over 250K timesteps with 5 seeds. Specifically, our framework loads all pre-trained networks, including the policy network, ensemble Q network and ensemble target Q network, while appending the necessary temperature hyperparameter for SAC to facilitate further fine-tuning. In the Antmaze environment, we maintain the same reward modification as the offline phase during training but keep the rewards unchanged during evaluation. To ensure a fair comparison with IQL and PEX, which utilize offline data, we also load LAPO and ENOTO-LAPO with offline data for online fine-tuning.
@@ Appendix D Additional Results
In this section, we provide more experiments and detailed results to help understand our proposed ENOTO framework more comprehensively.
@@ D.1 Ablation on Offline Data
We conduct an ablation study to investigate the impact of using offline data during the online fine-tuning phase for all MuJoCo locomotion tasks, as shown in Fig. 7. Our results show that ENOTO-CQL_buffer, which initializes the online buffer with offline data, exhibits slow performance improvement, while discarding the offline data allows it to achieve higher sample efficiency. This suggests that although many offline-to-online RL methods utilize offline data to alleviate performance degradation, it can adversely affect their online sample efficiency. In contrast, ENOTO-CQL successfully avoids significant performance drop even without using offline data, thereby enhancing learning efficiency during the online stage.
D.2 Comparison of LAPO and ENOTO-LAPO
Figure 8: Online learning curves of LAPO and ENOTO-LAPO across five seeds on Antmaze tasks. The solid lines and shaded regions represent mean and standard deviation, respectively.
On the Antmaze tasks, we have conducted a comparative analysis between our ENOTO-LAPO method and several offline-to-online RL methods. To further validate the effectiveness of our ENOTO framework, we directly utilize LAPO with two Q networks for offline-to-online training and compare it with our ENOTO-LAPO method. The results are shown in Fig. 8. As original LAPO can achieve near-optimal performance in simple environments such as umaze and medium mazes during the offline stage, the online fine-tuning performance of both LAPO and ENOTO-LAPO is comparable. However, in the more challenging large maze environment, directly using the offline pre-trained LAPO agent for online fine-tuning leads to slow performance improvement. By employing our proposed ENOTO framework, we demonstrate that ENOTO-LAPO can not only enhance the offline performance of LAPO, but also facilitate more rapid performance improvement while maintaining the offline performance without degradation. This approach enables the offline agent to quickly adapt to the real-world environment, providing efficient and effective online fine-tuning.
@@ D.3 Visualization and Analysis
Figure 9: Visualization of the distribution of states generated by ENOTO-CQL, CQL in online phase and states in offline dataset.
To better understand the training efficiency of ENOTO in comparison to traditional pessimistic offline RL algorithms, we compare the distribution of states generated by CQL, ENOTO-CQL in the online phase, and the distribution of states from the offline dataset. To visualize the results clearly, we plot the distribution with t-Distributed Stochastic Neighbor Embedding (t-SNE) Hinton and Roweis [2002]. The results are shown in Fig. 9, it can be found that both the distribution of ENOTO-CQL states and CQL states bear some similarity to the distribution of offline states. However, the online states accessed by CQL are located on the edge of the offline area, but most still overlap with the offline states. On the other hand, the online states accessed by ENOTO-CQL deviate further from the offline states. With our ensemble-based design for optimistic exploration, ENOTO empowers the offline agent to explore more states beyond those contained in the offline dataset. This capability allows for swift adaptation to online environments and facilitates rapid performance improvement.
D.4 Detailed Results of ENOTO Components
In this section, we provide all learning curves of ENOTO components that are restricted by the length of the text in the main paper.
@@ D.4 Detailed Results of ENOTO Components
Fig. 10 illustrates the performance of various offline-to-online RL approaches on MuJoCo locomotion tasks. It is evident that the OfflineRL $\to$ OnlineRL method exhibits the best performance in the HalfCheetah environment. However, it demonstrates unstable learning in the more complex environments of Walker2d and Hopper. On the other hand, the OfflineRL $\to$ OfflineRL approach remains stable but shows slower asymptotic performance. In contrast, the OfflineRL-N $\to$ OnlineRL-N method no longer experiences performance collapse after transitioning to online fine-tuning, and its training process is relatively stable across all tasks. Additionally, OfflineRL-N $\to$ OnlineRL-N achieves superior fine-tuned performance compared to OfflineRL $\to$ OfflineRL. It is worth noting that the Hopper-medium-expert-v2 dataset represents a special case where most considered offline-to-online RL methods exhibit varying degrees of performance drop, as depicted in this figure and subsequent figures. Nevertheless, our ENOTO framework consistently achieves state-of-the-art performance in comparison to all baseline methods across most tasks.
@@
Fig. 11 displays the performance of OnlineRL-N utilizing different Q-target computation methods on MuJoCo locomotion tasks. It is evident that MinQ exhibits remarkable stability across all tasks, albeit with slower performance improvement in the HalfCheetah and Walker2d environments. MeanQ and REM demonstrate excellent performance in the HalfCheetah environment, but struggle to improve in the more challenging environments of Walker2d and Hopper, and their learning process is characterized by instability. In contrast, RandomMinPair and WeightedMinPair showcase superior performance across most tasks, with the exception of the Hopper-medium-replay-v2 dataset where they exhibit slight instability in learning. Among these two methods, WeightedMinPair demonstrates slightly better stability and performance, thus we select it as the component of our final ENOTO framework and present the experiments related to RandomMinPair in the appendix.
@@
Fig. 12 and Fig. 13 present the performance of OnlineRL-N + WeightedMinPair and OnlineRL-N + RandomMinPair using different exploration methods on MuJoCo locomotion tasks, respectively. These two figures exhibit similar observations. In the Hopper environment, OAC achieves the best performance, but its performance improvement in HalfCheetah and Walker2d is relatively slow. The use of Bootstrapped DQN leads to minimal improvement in performance, while SUNRISE enhances the learning efficiency of OnlineRL-N + WeightedMinPair across most tasks, with the exception of the Hopper-medium-replay-v2 dataset where they exhibit slight instability in learning.
@@ D.5 Offline Performance
Fig. 14 shows the offline performance of different methods on MuJoCo locomotion tasks. We observe that on certain datasets such as HalfCheetah-medium-v2 and HalfCheetah-medium-replay-v2, ENOTO-CQL exhibits a slight performance improvement compared to CQL. This indicates that the introduction of Q-ensembles indeed has some benefits for the performance in the offline stage. However, it is important to note that the use of Q-ensembles can impact the convergence speed. For instance, on the Hopper-medium-expert-v2 dataset, ENOTO-CQL demonstrates a noticeably slower convergence speed compared to CQL. Nevertheless, if both algorithms are allowed to continue training, for example, for 3M training steps, we believe that the performance of ENOTO-CQL can still surpass that of CQL.
Although different offline RL algorithms may have varying final performance in the offline stage, with some methods potentially performing worse than ENOTO-CQL, our comparisons in the online stage remain fair. On one hand, the introduction of Q-ensembles can enhance the performance of existing algorithms in the offline stage, which is a inherent advantage of Q-ensembles. On the other hand, higher performance in the offline stage can actually lead to performance drop in the online stage, while lower performance in the offline stage is less prone to such drop. Our proposed ENOTO framework ensures that the offline policy maintains high performance from the offline stage, and achieves rapid performance improvement in the online stage without encountering performance drop.
@@
Fig. 15 displays the offline performance of various methods on Antmaze navigation tasks. Firstly, we observe that LAPO outperforms IQL in terms of offline performance, providing a higher starting point for the online phase. This is particularly evident in the umaze and medium maze environments, where LAPO nearly reaches the performance ceiling. Regarding LAPO and ENOTO-LAPO, since LAPO achieves near-optimal performance in simple environments such as umaze and medium mazes, their offline performance is comparable. However, in the more challenging large maze environment, the inclusion of Q-ensembles enables ENOTO-LAPO to surpass LAPO in terms of performance.